Example of the Use of MSLAPA Bibliography Style

Michael S. Landy

1 Introduction

blah blah blah (Aloimonos & Shulman, 1989, note1). blah blah blah.

Another paragraph of blah blah blah (note2, Braunstein, 1968; Aloimonos & Shulman, 1989). A nocite-ation of Bergen and Landy ’91.

1.1 Stuff I understand

blah blah blah.

1.2 Stuff I made up

blah blah (note3, Bergen & Landy, 1991, note4) blah. blah blah (Dosher, Landy & Sperling, 1989a,b; Dosher, Sperling & Landy, 1989c). blah blah in the work of Dosher and colleagues (1989a). blah blah in the work of Dosher and colleagues (1989b,c). blah blah in the work of Dosher and colleagues (1989b, 1990). blah blah (Dosher et al., 1989a). blah blah (Aloimonos & Shulman, 1989; Dosher et al., 1989a,b).

References

Aloimonos, J. & Shulman, D. (1989). Integration of Visual Modules. New York: Academic Press.

Bergen, J. R. & Landy, M. S. (1991). Computational modeling of visual texture segregation. In M. S. Landy & J. A. Movshon (Eds.), Computational Models of Visual Processing (pp. 253–271). Cambridge, MA: MIT Press.

Braunstein, M. L. (1968). Motion and texture as sources of slant information. Journal of Experimental Psychology, 78, 247–253.

Dosher, B. A., Landy, M. S. & Sperling, G. (1989a). The kinetic depth effect and optic flow — I. 3D shape from fourier motion. Vision Research, 29, 1789–1813.

Dosher, B. A., Landy, M. S. & Sperling, G. (1989b). Ratings of kinetic depth in multi-dot displays. Journal of Experimental Psychology: Human Perception and Performance, 15, 816–825.

Dosher, B. A., Sperling, G. & Landy, M. S. (1989c). The kinetic depth — A nonexistent article. Vision Research, 29, 100–113.

Dosher, B. A., Sperling, G. & Landy, M. S. (1990). The kinetic depth — Another nonexistent article. Vision Research, 29, 100–113.